# On functoriality and the tensor product property in noncommutative tensor-triangular geometry

### v2

@@ 8. The complete prime spectra of crossed product categories
Set $\mathscr{K}=D_{b}(k)\oplus D_{b}(k)$. The crossed product category $\mathscr{K}\rtimes C_{2}$, with $C_{2}$ acting by permuting indices, contains no completely prime ideals. Moreover, the inclusion $F:\mathscr{K}\hookrightarrow\mathscr{K}\rtimes C_{2}$ does not induce a well-defined map $\operatorname{Spc}(F):\operatorname{Spc}(\mathscr{K}\rtimes C_{2})\to\operatorname{Spc}(\mathscr{K})$ nor a well-defined functor $\mathscr{S}(F):\mathscr{S}(\mathscr{K}\rtimes C_{2})\to\mathscr{S}(\mathscr{K})$.
@@
Next, consider the inclusion $F:\mathscr{K}\hookrightarrow\mathscr{K}\rtimes C_{2}$ defined by $x\mapsto x\boxtimes 1_{G}$. This is a monoidal triangulated functor. We have that $F^{-1}(\{0\})=\{0\}$, which is a nonprime thick $\otimes$-ideal, so the map $F^{-1}$ does not send primes to primes.
@@ 9. Duality and duo-ness
$(\operatorname{ev}_{x}^{\prime}\otimes\operatorname{id}_{x})\circ(\operatorname{id}_{x}\otimes\operatorname{coev}_{x}^{\prime})=\operatorname{id}\text{ and }(\operatorname{id}_{{}^{\vee}x}\otimes\operatorname{ev}_{x}^{\prime})\circ(\operatorname{coev}^{\prime}_{x}\otimes\operatorname{id}_{{}^{\vee}x})$
@@
First we show (a) implies (b) and (c). Suppose every thick right $\otimes$-ideal of $\mathscr{K}$ is two-sided. It suffices to show that if $x\in\mathscr{I}$, then $x^{\vee}\in\mathscr{I}$ and ${}^{\vee}x\in\mathscr{I}$, the other implications follow by replacing $x$ with $x^{\vee}$ or ${}^{\vee}x$. Since $\mathscr{I}$ is two-sided, $x^{\vee}\otimes x\otimes x^{\vee}\in\mathscr{I}$. Because $\mathscr{I}$ is thick and $x^{\vee}$ is a summand of $x^{\vee}\otimes x\otimes x^{\vee}$, $x^{\vee}\in\mathscr{I}$. The right dual case follows similarly. (b) implies (c) is obvious.
@@ 10. Surjectivity and tensor-nilpotence
:=$\mathscr{I}^{\prime}:=\{y\in\mathscr{L}\mid\exists\,x\in\mathscr{P}\text{ suchthat }y\in\langle F(x)\rangle_{r}\}.$
@@
There are two key points which prevent us from fully extending [Bal18, Theorem 1.4], a stronger statement that characterizes exactly when $\varphi:\operatorname{Spc}(\mathscr{L})\to\operatorname{Spc}(\mathscr{K})$ is surjective in generality. In one direction (the direction (a) $\implies$ (b) in the proof), it requires that the collection $\{z\in\mathscr{K}\mid f\text{ is $\otimes$-nilpotent on }z\}$ be a thick triangulated $\otimes$-ideal, which as noted in 10.2, may not be a triangulated subcategory. Additionally, in the (b) $\implies$ (a) direction, it is deduced from the fact that $\xi_{x}^{\otimes n}$ is $\otimes$-nilpotent on its cone that $g=\xi_{x}^{\otimes n}\otimes s$ is $\otimes$-nilpotent on its cone as well; this does not hold in our case, even if we replace $\otimes$-nilpotent” with “weakly $\otimes$-nilpotent.”
@@
[…]
@@ 11. Finale: the universal functorial support theory
A proper partial $\otimes$-ideal $\mathscr{I}$ is prime if furthermore, the following holds for all commuting $x,y\in\mathscr{K}$: if $x\otimes y\in\mathscr{I}$, then $x\in\mathscr{I}$ or $y\in\mathscr{I}$. In particular, one has that for any pair of commuting $x,y\in\mathscr{K}$, $x\otimes y\in\mathscr{I}$ if and only if $x\in\mathscr{I}$ or $y\in\mathscr{I}$.
@@
Let $F:\mathscr{K}\to\mathscr{L}$ be an exact $\otimes$-functor. If $\mathscr{I}$ is a partial $\otimes$-ideal of $\mathscr{L}$, then $F^{-1}(\mathscr{I})$ is a partial $\otimes$-ideal of $\mathscr{K}$. Moreover, if $\mathscr{I}$ is prime, then so is $F^{-1}(\mathscr{I})$.
@@
Take the subcategory of $\mathscr{K}$ given by finite coproducts and tensor products of $x$ and $y$ containing only isomorphisms and zero maps for morphisms. This category is trivially triangulated and monoidal, and a braiding can be endowed by choosing an isomorphism $x\otimes y\cong y\otimes x$ and extending it to all products of $x$ and $y$ this satisfies the hexagon identities by construction. ∎
@@
Let $F:\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Top}$ be a functor for which there exists a natural transformation $\eta:F\to\operatorname{Spc}$ upon restriction to $\mathbf{Bmon}_{\Delta}$. We must show there is a unique natural transformation $\delta:F\to p\operatorname{Spc}$ that when restricted to $\mathbf{Bmon}_{\Delta}$ induces $\eta$. For every essentially small monoidal triangulated category $\mathscr{K}$ and every braided monoidal triangulated subcategory $\mathscr{C}\subseteq\mathscr{K}$, the inclusion induces a continuous map $F(\mathscr{K})\to F(\mathscr{C})$, and we have a morphism $\eta_{\mathscr{C}}:F(\mathscr{C})\to\operatorname{Spc}(\mathscr{C})$, hence a morphism $F(\mathscr{K})\to\operatorname{Spc}(\mathscr{C})$. Therefore, we have a cone over the diagram $\operatorname{Br}(\mathscr{K})^{\operatorname{op}}$, so by universality of the inverse limit, we have a unique morphism
@@
The functor $p\operatorname{Spc}:\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Top}$ is the final object in the full subcategory of $\operatorname{Fun}(\mathbf{mon}_{\Delta}^{\operatorname{op}},\mathbf{Top})$ consisting of functors $F:\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Top}$ which have a natural transformation to $\operatorname{Spc}$ upon restriction to $\mathbf{Bmon}_{\Delta}^{\operatorname{op}}$. In particular, it is the final object in the full subcategory consisting of functors whose natural transformations are natural isomorphisms as well, i.e. the fiber category $\iota^{-1}(\operatorname{Spc})$, where $\iota$ denotes the inclusion $\mathbf{Bmon}_{\Delta}^{\operatorname{op}}\hookrightarrow\mathbf{mon}_{\Delta}^{\operatorname{op}}$.

### v3

@@ 8. The complete prime spectra of crossed product categories
Set $\mathscr{K}=D_{b}(k)\oplus D_{b}(k)$. The crossed product category $\mathscr{K}\rtimes C_{2}$, with $C_{2}$ acting by permuting indices, contains no completely prime ideals. Moreover, the inclusion $F\colon\mathscr{K}\hookrightarrow\mathscr{K}\rtimes C_{2}$ does not induce a well-defined map $\operatorname{Spc}(F):\operatorname{Spc}(\mathscr{K}\rtimes C_{2})\to\operatorname{Spc}(\mathscr{K})$ nor a well-defined functor $\mathscr{S}(F):\mathscr{S}(\mathscr{K}\rtimes C_{2})\to\mathscr{S}(\mathscr{K})$.
@@
Next, consider the inclusion $F\colon\mathscr{K}\hookrightarrow\mathscr{K}\rtimes C_{2}$ defined by $x\mapsto x\boxtimes 1_{G}$. This is a monoidal triangulated functor. We have that $F^{-1}(\{0\})=\{0\}$, which is a nonprime thick $\otimes$-ideal, so the map $F^{-1}$ does not send primes to primes.
@@ 9. Duality and duo-ness
[…]
@@
First we show (a) implies (b) and (c). Suppose every thick right $\otimes$-ideal of $\mathscr{K}$ is two-sided. It suffices to show that if $x\in\mathscr{I}$, then $x^{\vee}\in\mathscr{I}$ and $^{\vee}x\in\mathscr{I}$, the other implications follow by replacing $x$ with $x^{\vee}$ or $^{\vee}x$. Since $\mathscr{I}$ is two-sided, $x^{\vee}\otimes x\otimes x^{\vee}\in\mathscr{I}$. Because $\mathscr{I}$ is thick and $x^{\vee}$ is a summand of $^{\vee}\otimes x\otimes x^{\vee}, xx^{\vee}\in\mathscr{I}$. The right dual case follows similarly. (b) implies (c) is obvious.
@@ 10. Surjectivity and tensor-nilpotence
$\mathscr{I}^{\prime}:=\{y\in\mathscr{L}\mid\exists\,x\in\mathscr{P}\text{ such that }y\in\langle F(x)\rangle_{r}\}.$
@@
[…]
@@
Now a morphism $f\colon x\to y$ satisfies $F(f)=0$ if and only the composition $x\xrightarrow{f}y\xrightarrow{u\otimes y}A\otimes y$ is zero. Indeed, this follows from the projection formula $A\otimes-\cong UF(-)$ with $y=\mathbbm{1}_{\mathscr{L}}$. Now, this is equivalent to the morphism $f\colon x\to y$ factoring via $\xi\otimes y\colon J\otimes y\to y$, via the exact triangle
@@ 11. Finale: the universal functorial support theory
A proper partial $\otimes$-ideal $\mathscr{I}$ is prime if furthermore, the following holds for all commuting $x,y\in\mathscr{K}$: if $x\otimes y\in\mathscr{I}$, then $x\in\mathscr{I}$ or $y\in\mathscr{I}$. In particular, one has that for any pair of commuting $\in\mathscr{K}, xyx\otimes y\in\mathscr{I}$ if and only if $x\in\mathscr{I}$ or $y\in\mathscr{I}$.
@@
Let $F\colon\mathscr{K}\to\mathscr{L}$ be an exact $\otimes$-functor. If $\mathscr{I}$ is a partial $\otimes$-ideal of $\mathscr{L}$, then $F^{-1}(\mathscr{I})$ is a partial $\otimes$-ideal of $\mathscr{K}$. Moreover, if $\mathscr{I}$ is prime, then so is $F^{-1}(\mathscr{I})$.
@@
Take the subcategory of $\mathscr{K}$ given by finite coproducts and tensor products of $x$ and $y$ containing only isomorphisms and zero maps for morphisms. This category is trivially triangulated and monoidal, and a braiding can be endowed by choosing an isomorphism $x\otimes y\cong y\otimes x$ and extending it to all products of $x$ and $y$ - this satisfies the hexagon identities by construction. ∎
@@
Let $F\colon\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Top}$ be a functor for which there exists a natural transformation $\eta:F\to\operatorname{Spc}$ upon restriction to $\mathbf{Bmon}_{\Delta}$. We must show there is a unique natural transformation $\delta:F\to p\operatorname{Spc}$ that when restricted to $\mathbf{Bmon}_{\Delta}$ induces $\eta$. For every essentially small monoidal triangulated category $\mathscr{K}$ and every braided monoidal triangulated subcategory $\mathscr{C}\subseteq\mathscr{K}$, the inclusion induces a continuous map $F(\mathscr{K})\to F(\mathscr{C})$, and we have a morphism $\eta_{\mathscr{C}}:F(\mathscr{C})\to\operatorname{Spc}(\mathscr{C})$, hence a morphism $F(\mathscr{K})\to\operatorname{Spc}(\mathscr{C})$. Therefore, we have a cone over the diagram $\operatorname{Br}(\mathscr{K})^{\operatorname{op}}$, so by universality of the inverse limit, we have a unique morphism
@@
The functor $p\operatorname{Spc}:\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Top}$ is the final object in the full subcategory of $\operatorname{Fun}(\mathbf{mon}_{\Delta}^{\operatorname{op}},\mathbf{Top})$ consisting of functors $F\colon\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Top}$ which have a natural transformation to $\operatorname{Spc}$ upon restriction to $\mathbf{Bmon}_{\Delta}^{\operatorname{op}}$. In particular, it is the final object in the full subcategory consisting of functors whose natural transformations are natural isomorphisms as well, i.e. the fiber category $\iota^{-1}(\operatorname{Spc})$, where $\iota$ denotes the inclusion $\mathbf{Bmon}_{\Delta}^{\operatorname{op}}\hookrightarrow\mathbf{mon}_{\Delta}^{\operatorname{op}}$.